OpenStudy (anonymous):

Help with a geometry problem, triangles. In a triangle ABC, it draws the interior line BD, for which AD=BC , if the angle A is 54°, the angle C is 24° y the angle DBC is x, find x. We are supposed to find some congruent triangles here. Any idea.?

OpenStudy (anonymous):

Lets draw the triangle first.

OpenStudy (anonymous):

does it give you a figure already?

OpenStudy (anonymous):

Here is the graph.

OpenStudy (anonymous):

okay give me a minute to open up the file :)

OpenStudy (anonymous):

The answer is E. 18degrees

OpenStudy (anonymous):

Let me draw you a graph

OpenStudy (anonymous):



OpenStudy (anonymous):

this is just a way to look at the question and of course youhave many ways of solving this question.

OpenStudy (anonymous):

so you know that the line where the 54, y, x and 24 degrees become a 180 degree supplementary line. so we do 180-54-24=120, so we know that x+y=102

OpenStudy (anonymous):

then you look again and see oh x and y are not complementary angels. that means they must be smaller than 90 degrees. a biggest not angel not equal to 90 degrees is 89. so 102-89=13, lets start from the closest #, 102-10=92, wrong, its bigger than 90 degrees, 102-15=93, still wrong, 102-18=84! Bingo, thats the one.

OpenStudy (anonymous):

on the last post I meant 180-54-24=102

OpenStudy (anonymous):

thats the elimanate watyy to do it.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Well, it´s clear for me that x+y=102, but where does come from that they both must be less than 90?

OpenStudy (anonymous):

because using the triangel identities that if BDU is a obese angle then that means there cannot be any other angle bigger than 90 degrees. and if n a nother trianglethere are two angles together added up more than 90 degrees the third angle is not a90 degree angle.

OpenStudy (anonymous):

@jokereq7 The problem don´t say anithing about obtuse triangles, all the information given is the measures of those angles and that AD = BC. And if this is this reasoning is correct, why they say that AD=BC? Shouldn´t that information be used in order to solve the problem? By this reasoning the answer could be any angle above 12°, why not 15°? What do you think?

OpenStudy (anonymous):

Answer is 10

OpenStudy (anonymous):

@bhaskarbabu Can you show us how did you get that result?

OpenStudy (anonymous):

in the same way as we did the day before yesterday

OpenStudy (anonymous):

But here the triangle ABD is not isosceles. Its altitude is not median. How we could relate the sides AD and BC drawing an altitude?

OpenStudy (anonymous):

first draw an altitude and find how it divides AD then continue as the previous one

OpenStudy (anonymous):

if AD= a then assume the parts of AD as 'a' and 'a-y' then find y

OpenStudy (anonymous):

 So i get a sin(24) = (a-y)tan(54) And tan (24+x) = [asin(24)]/y Well, solve y in terms of a may take a while without a calculator. Doing with a calculator i get that the a factor cancels and i get. tan( 24 + x) = 0.58 And x=10 don´t satisfies that. Where i (or the calculator) went wrong ?